A force F is pulling a 16.0 kg box along a 30 Degree  ramp at a constant velocity. Note that the force F is acting along the ramp surface. Kinetic fictional coefficient of ramp surface is 0.57. Calculate the work done by force F to move the box 8.0 m along that surface? (Hint: Find F first). g=10 m/s^2
Answer #1

Since it is with constant velocity accelration would be 0.

N=mgcos\theta and F= \muNsin\theta

so F=\mumg cos\thetasin\theta

Work done by force = F*d = \mumg cos\thetasin\theta*d

putting in values we get Work done = .57 *16*10*(1.732/4)* 8=315.92 J